Find a power series solution of xy'=y​

1
Expert's answer
2022-01-30T15:08:48-0500

Let

"y\left(x\right)=\sum_{n=0}^{\infty}{a_nx^n}"

"\frac{dy}{dx}=\sum_{n=1}^{\infty}{na_nx^{n-1}}=\sum_{n=0}^{\infty}{\left(n+1\right)a_{n+1}x^n}"

Substitution into equation:

"\sum_{n=0}^{\infty}{\left(n+1\right)a_{n+1}x^{n+1}}=\sum_{n=0}^{\infty}{a_nx^n}"

"\sum_{n=1}^{\infty}{na_nx^n}-\sum_{n=0}^{\infty}{a_nx^n}=0"

"-a_0+\sum_{n=1}^{\infty}{na_nx^n}-\sum_{n=1}^{\infty}{a_nx^n}=0"

Coefficients near xn are equal to zero.

n=0: a0=0  

n=1: a1 - a1=0   =>  a1 is arbitrary constant C

n>1: nan - an=0  =>  (n-1)an=0  =>  an = 0 

Therefore for arbitrary a1 = C and an = 0 for any n≠1 we will get solution y(x) = Cx

Answer

y(x) = Cx